OpenStudy (anonymous):

Prove how to get future value formula. FV=p(1+r/m)^mn

OpenStudy (anonymous):

Break the yearly rate r into its monthly rate r/m There are nm such months in total After first month FV(1) = P (1+r/m) After second month FV(2) = FV(1) (1+r/m) or FV(2) = P (1+r/m)^2 and so on at the end of total period, mn months FV(mn) = P (1+r/m)^mn